OpenStudy (anonymous):

Can someone please help me with this one. Or just explain it to me. Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. A. y = -6x - 15 B. y = -6x - 20 C. y = -6x + 15 D. y = -6x + 20

OpenStudy (anonymous):

you are given an ordered pair (x,y) in (-2,-3). just plug those values into each of those equations(A through D) and see which equation is satisfied by the ordered pair you have been given.

OpenStudy (anonymous):

Still confused

OpenStudy (anonymous):

A

OpenStudy (anonymous):

I'll show you how to do D and then you can do the rest. x = -2 y = -3 y = -6x + 20 -3 = (-6)(-2) + 20 -3 = 12 + 20 -3 = 32 obviously this is not the right equation. that should clear things up.

OpenStudy (anonymous):

@Sky_Pirate now you have the responsibility to explaining how you found that answer. don't just give it away without explaining.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@tomo

OpenStudy (anonymous):

how did you get 20?

OpenStudy (anonymous):

its given in the initial equation for D.

OpenStudy (anonymous):

i still dont get it haha its okay, ill try to figure it out. but thanks though